# DiaSynth - Synthetic Dialogue Generation Framework

### v1

@@ 5.1 Quality of the Dialogues
Table 5: G_EVAL
5.2 Fine-tuning and Performance Results
The experimental setup is designed as follows:
  • All reported metrics are based on the validation and test sets of DialogSum and SAMSum to maintain consistency in evaluation.
  • To assess the effectiveness of DiaSynth-generated data, we compare the performance of models fine-tuned on DiaSynth data with their corresponding base models, which are evaluated without any fine-tuning.
  • Given that the training sets of DialogSum and SAMSum are larger than the synthetic datasets generated in this study, we randomly sampled instances from the in-domain training sets to match the size of the DiaSynth-generated data. This ensures a fair comparison in terms of training scale.
  • The primary goal of the experiment is to quantify the percentage improvement offered by models fine-tuned on DiaSynth-generated data compared to their base versions and to evaluate the extent to which the generated data capture the distributional characteristics of the in-domain data.
The results presented in Table 6 highlight the performance of each model before and after fine-tuning. Notably, base models such as DistilBART and LED exhibit relatively high QAGS scores before fine-tuning, primarily because these models tend to repetitively generate content similar to the input dialogues, thereby artificially inflating the factual consistency metric. This repetition, while beneficial for preserving factual alignment, does not necessarily translate into improved quality or informativeness of the summaries.
After fine-tuning on DiaSynth-generated data, models show a significant increase in BERTScore and ROUGE-L compared to their base versions, demonstrating DiaSynth’s effectiveness in enhancing model performance for downstream tasks. This result suggests that DiaSynth is valuable for generating dialogue data in domains with limited training resources. Notably, models fine-tuned on GPT-4o-generated data performed best for DialogSum, while those fine-tuned on LLaMA-3-generated data excelled on SAMSum, further reinforcing that open-source LLMs are more adept at generating informal, casual dialogues like those in SAMSum. The percentage improvement of models fine-tuned on LLaMA-3-generated data was calculated using Equation 1, showing an average improvement of 16.47% across the 24 reported results. We chose to calculate on LLaMA-3 because across both quality and usability it performs better than GPT-4o.
Percentage Improvement=Score After Fine-tuningScore Before Fine-tuningScore Before $\text{Percentage Improvement}=\frac{\text{Score After Fine-tuning}-\text{ScoreBefore Fine-tuning}}{\text{Score Before Fine-tuning}}\times 100$ (1)
Coverage Percentage=Finetuned Score on DiaSynth DataFinetuned Score on In-Domain Base $\text{Coverage Percentage}=\frac{\text{Finetuned Score on DiaSynth Data}}{\text{Finetuned Score on In-Domain Base Data}}\times 100$ (2)
To assess the percentage coverage of the distributional characteristics of the in-domain data by the synthetically generated data, we use Equation 2 and the scores of models fine-tuned on LLaMA-3 generated data. Across the 24 reported results, the overall coverage percentage of the LLaMA-3 generated data is 90.48%. Notably, the QAGS scores of models fine-tuned on synthetic data surpass those of models trained on in-domain data, suggesting that synthetic data can match or even exceed in-domain data performance in some aspects. Excluding QAGS, the coverage percentage is calculated to be 77.07%, highlighting that even with moderate LLMs of small scale (3B - 8B), high-quality synthetic dialogue datasets can be effectively created.
Model Before Finetuning Finetuning on Base Data
@@ 5.2 Fine-tuning and Performance Results
Table 7: Performance metrics
6 Limitations
Despite the promising results, our approach has some limitations. Firstly, different LLMs exhibit varied performance based on the dialogue structure, with certain models like LLaMA-3 performing better for more informal dialogues (e.g., SAMSum) and others like GPT-4o excelling in structured, formal dialogues (e.g., DialogSum). This indicates that there is no single model that can universally handle all types of dialogue structures, requiring a tailored choice of LLM depending on the target application.
Secondly, the generation process may suffer from a lack of knowledge on certain topics, especially in cases where the LLMs were not sufficiently trained on those domains. Additionally, our framework relies on zero-shot generation for personas and sub-topics, which, while flexible, can sometimes result in less coherent or less accurate persona simulation, as it is not fine-tuned for specific contexts. These limitations suggest directions for future work, such as combining LLMs to leverage their strengths or incorporating more topic-specific training to enhance knowledge coverage.
@@ 7 Conclusion
Appendix A Scalability of DiaSynth

### v2

@@ 5.1.1 Metric Scores
Table 4: G-EVAL
5.1.2 Strong performance of LLaMA-3
Model Before Finetuning Finetuning on In-Domain Data
@@ 5.2 Fine-tuning and Performance Results
Table 6: Performance after finetuning on synthetic data
The experimental setup is designed as follows:
  • Metrics are reported on the validation and test sets of DialogSum and SAMSum.
  • To evaluate DiaSynth-generated data, we compared models fine-tuned on DiaSynth data with their base versions (no fine-tuning).
  • In-domain training sets were randomly sampled to match the size of the DiaSynth-generated data, enabling fair comparison.
  • The experiment aimed to quantify performance improvement of DiaSynth-fine-tuned models and assess alignment with in-domain data distributions.
  • Models were fine-tuned for 2 epochs with a learning rate of 5e-5 and a warmup of 50 steps.
The results presented in Tables 5 and 6 present the performance of the base models, models finetuned on in-domain data and models finetuned on DiaSynth generated data. Models finetuned on DiaSynth data generally improves the performances from the BERTScore and ROUGE-L metrics. Surprisingly, for some models (LED and BART) the QAGS scores were higher than the models finetuned on DiaSynth. On further exploration, we found out that these models extracted multiple sentences from the given dialogue instead of generating a summary which led to high QAGS scores. Comparing models finetuned on in-domain data to those finetuned on DiaSynth data reveals that DiaSynth finetuning generally enhances factual accuracy, with BERTScore and ROUGE-L scores remaining comparable. The disparity in BERTScore and ROUGE-L results may be due to format variations. Models fine-tuned on in-domain data were evaluated on summaries that matched the training format closely, while DiaSynth-fine-tuned models were trained on LLM-generated summaries and evaluated on human-generated summaries, leading to minor format mismatches. Comparison between the different LLMs from Table 6, shows that GPT-4o is better at generating dialogues and summaries that are formal in nature while LLaMA-3 and open source LLMs would be better for generating dialogues that are informal and casual in nature.
% Improvement $=\frac{\text{After - Before}}{\text{Before}}$ (1a)
% Coverage $=\frac{\text{Score DiaSynth}}{\text{Score In-domain}}$ (1b)
Model % Improvement % Covered
distilbart-cnn 10.96 88.81
bart-base 9.21 90.6
t5-base 7.59 93.67
led-base-16384 2.14 89.68
Table 7: Summarization results on DialogSum
Model % Improvement % Covered
distilbart-cnn 6.07 87.25
bart-base 16.12 94.35
t5-base 30.04 87.36
led-base-16384 15.25 90.91
Table 8: Summarization results on SAMSum
To assess the percentage improvement and percentage coverage of the distributional characteristics of the in-domain data by the synthetically generated data, we use Equations 1a and 1b respectively. We use the scores of models finetuned on LLaMA-3 generated data because of its dominance in both quality and usability. Across the 24 reported results, the overall coverage percentage of the LLaMA-3 generated data is 90.48%. Notably, the QAGS scores of models fine-tuned on synthetic data surpass those of models trained on in-domain data, suggesting that synthetic data can match or even exceed in-domain data performance in some aspects. Excluding QAGS, the coverage percentage is calculated to be 77.07%. In addition to the average percentages, we also present the model wise percentage improvement and coverage in Table 8 and 8. The results presented are with respect to the dialogues generated using LLaMA-3 and they illustrate clear improvements for every model, highlighting that even with moderate LLMs of small scale (3B - 8B), high-quality synthetic dialogue datasets can be effectively created across different domains and different dialogue formats.
@@ 6 Conclusion
7 Limitations
Despite the promising results, our approach has some limitations. Firstly, different LLMs exhibit varied performance based on the dialogue structure, with certain models like LLaMA-3 performing better for more informal dialogues (e.g., SAMSum) and others like GPT-4o excelling in structured, formal dialogues (e.g., DialogSum). This indicates that there is no single model that can universally handle all types of dialogue structures, but a single SOTA model can give stable and decent results like LLaMA-3.
Secondly, the generation process may suffer from a lack of knowledge on certain topics, especially in cases where the LLMs were not sufficiently trained on those domains. Additionally, our framework relies on zero-shot generation for personas and sub-topics, which, while flexible, can sometimes result in less coherent or less accurate persona simulation, as it is not fine-tuned for specific contexts.
Since LLMs power DiaSynth, hallucinations and compute-need are two inherent limitations. We present a detailed hallucination study in Appendix A, which indicates that though the hallucinations rates are acceptable, there is still scope for improvements. These limitations suggest directions for future work, such as combining LLMs to leverage their strengths or incorporating more topic-specific training to enhance knowledge coverage.
Appendix A Hallucination Study
In addition to evaluating the quality and usability of dialogues produced by DiaSynth, we conducted a study on the phenomenon of hallucinations within the generated dialogues. Hallucinations in language models refer to instances where the output contains misleading or incorrect information or situations where the model repeats the same content. To evaluate the occurrence of hallucinations, we compared the generated dialogues with their respective summaries and assessed them using two well-known hallucination benchmarks: SelfCheckGPT Manakul et al. [2023] and ChainPoll Friel and Sanyal [2023]. This analysis provides insights into the prevalence of hallucinations and informs strategies for improving dialogue quality in future iterations of DiaSynth. The results are presented in Tables 10 and 10.
A.1 SelfCheckGPT
SelfCheckGPT quantifies the self-consistency of LLM outputs by examining agreement across multiple outputs from the same prompt. This assessment reveals potential inaccuracies through metrics like SelfCheck-BertScore.
The SelfCheck-BERTScore results for various models show that hallucination levels are at worst around 25%, which is acceptable but still indicates areas for improvement. Across both datasets, Phi-3 demonstrates the most robustness, likely due to its pretraining on structured, textbook-like data, which may contribute to greater consistency and factual accuracy.
A.2 ChainPoll
ChainPoll utilizes a chain-of-thought prompting approach to identify hallucinations by iteratively polling the model with structured reasoning prompts. This method systematically detects both open-domain and closed-domain hallucinations, where lower scores indicate fewer hallucinations.
The ChainPoll scores indicate that hallucination levels on these models are generally low, with the best performance seen by GPT-4o on SAMSum, which achieves the lowest score of 0.120, suggesting minimal hallucinations. On the other hand, LLaMA-3 scores higher at 0.237 on SAMSum, indicating more frequent hallucinations. These findings highlight different models’ strengths in generating accurate and reliable dialogues.
A.3 Implications for DiaSynth
The results from both SelfCheckGPT and ChainPoll evaluations suggest that DiaSynth, when leveraging models like Llama 3, is capable of generating dialogues with relatively low hallucination rates. However, specific models show variability in performance across datasets, indicating that further enhancements, such as fine-tuning or incorporating additional guardrails, could improve DiaSynth’s robustness in generating reliable dialogues across diverse domains.
LLM ChainPoll SCGPT-BERTScore
Phi-3 0.198 0.791
InternLM-2.5 0.199 0.726
LLaMA-3 0.205 0.793
GPT-4o 0.178 0.765
Table 9: Hallucination calculation for DialogSum few-shot data
LLM ChainPoll SCGPT-BERTScore
Phi-3 0.154 0.785
InternLM-2.5 0.159 0.716
LLaMA-3 0.237 0.733
GPT-4o 0.120 0.742
Table 10: Hallucination calculation for SAMSum few-shot data
Appendix B Scalability of DiaSynth
@@ B.4 Scaling Observations
Appendix C Characteristics for the conversation
Table 11 shows different characteristics that we let the LLMs reason and decide using CoT. Before generating the dialogues, the LLMs are prompted to first reason about the various characteristics list for the dialogue given the topic and the personas and how the LLMs reason are illustrated in Appendix D.
Characteristic Description
Age and Gender Defines demographic details, influencing style and tone.
Familiarity Level Affects formality and depth based on relationship between speakers.
Emotional States Impacts tone and flow based on emotions (e.g., happy, sad).
Formality Level Determines level of politeness or casualness.
Duration of the Conversation Suggests the intended length and complexity of dialogue.
Communication Medium Defines the medium (e.g., face-to-face, phone), influencing style.
Topic of the Conversation Guides the content and direction of the dialogue.
Location of the Conversation Adds context influencing formality and content.
Agreement or Disagreement Drives dialogue dynamics based on agreement level.
Natural Dialogue Features Adds authenticity with fillers, pauses, and slang.
Table 11: Characteristics of the Dialogue for CoT Prompt
Appendix D Example CoT environments
These examples illustrate how CoT sets the various dialogue characteristics defined in Appendix C. As can be seen in the examples, the characteristics for the conversation are completely changed based on the personas and the topics, resulting in more grounded conversation generation. Future works can further explore how CoT can be used to further break down to generate even more realistic dialogues.
(a) CoT Example 1
(b) CoT Example 2
Figure 2: Illustration of Example CoT environments
Appendix E Ablation study on the use of CoT
In this section of the appendix, we present the impact of CoT in the dialogues generated by DiaSynth. We generate two datasets without CoT, using Phi-3 using DialogSum and SAMSum as few-shot examples with 8 topics. Tables 13, 13, 15 and 15 compare the scores of dialogues generated with and without CoT for the FED score and GPTScore and it can be clearly that CoT tends to increase the quality of the dialogues generated by DiaSynth.
We hypothesize that this improvement in quality is due to allowing the LLM to set diverse characteristics for the dialogue before generating the dialogue. This illustrates that either manually setting the relevant context or letting the LLM on its own to set the relevant context, we get better outputs, as adding relevant context lowers the probabilities of sequences that are not useful. The lower FED scores of CoT generated dialogues in Table 15, might be because of the CoT generated dialogues being longer in length but it needs further research.
Criteria Without CoT With CoT
Coherent 0.9507 0.9521
Error Recovery 0.938 0.9424
Consistent 0.9424 0.9523
Diverse 0.9431 0.952
Depth 0.9451 0.9506
Likeable 0.0088 -0.0003
Understand 0.9317 0.9338
Flexible -0.0013 0.0001
Informative 0.0032 0.0009
Inquisitive 0.0095 -0.003
Table 12: FED scores of dialogues generated with and without CoT for DialogSum few-shot
Criteria Without CoT With CoT
Coherence 0.0052 0.0284
Diversity 0.0120 0.0303
Flexibility 0.0074 0.0215
Understandability 0.0056 0.019
Inquistiveness 0.0162 0.0362
Consistency 0.0091 0.0366
Informativeness 0.0099 0.0168
Likeability 0.0029 0.0209
Depth 0.0062 0.0115
Error Recovery 0.0142 0.0242
Table 13: GPTScore of dialogues generated with and without CoT for DialogSum few-shot
Criteria Without CoT With CoT
Coherent 0.9667 0.9125
Error Recovery 0.9577 0.9051
Consistent 0.9621 0.9163
Diverse 0.9586 0.9124
Depth 0.9589 0.9094
Likeable 0.0059 -0.0003
Understand 0.9503 0.8978
Flexible -0.0022 -0.0023
Informative 0.0110 0.0035
Inquisitive 0.0030 -0.0037
Table 14: FED scores of dialogues generated with and without CoT for SAMSum few-shot
Criteria Without CoT With CoT
Coherence 0.0076 0.0324
Diversity 0.0172 0.0372
Flexibility 0.0116 0.0259
Understandability 0.0104 0.0272
Inquistiveness 0.0201 0.0389
Consistency 0.0149 0.0415
Informativeness 0.0148 0.0200
Likeability 0.0041 0.023
Depth 0.0078 0.0124
Error Recovery 0.0163 0.0289
Table 15: GPTScore of dialogues generated with and without CoT for SAMSum few-shot